40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 13 21:04:37 2013, Jenne, Update, Green Locking, PSL green cleaned up 
    Reply  Thu Mar 14 04:20:54 2013, Jenne, Update, Green Locking, Ybeat attempt 
       Reply  Thu Mar 14 20:22:33 2013, Jenne, Update, Green Locking, Xbeat attempt 
    Reply  Thu Mar 14 13:38:29 2013, Jenne, Update, Green Locking, PSL green cleaned up 
Message ID: 8297     Entry time: Thu Mar 14 20:22:33 2013     In reply to: 8291
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: Xbeat attempt 

I aligned the Xgreen and PSL green to overlap on the X beat PD, and reconnected the splitter which combines the X and Y beat signals and sends them to the control room.

I've been stepping the Xend laser temperature offset in steps of 20 counts, making sure the cavity unlocks and relocks on TEM00.  So far I have not seen any beat signals for the Xarm.  I've gone from 0 to 840.

I'll be back in a few hours to keep trying, although interested parties are invited to give it a whirl. 

 

ELOG V3.1.3-